# Having Second Thoughts? Let’s hear it

### v1

@@ 1 Introduction
Deep learning (DL) has been proven in various domains including computer vision and natural language processing [1, 2]. As deep neural networks or DL models rely on a large amount of linear/nonlinear operations, their operations are intractable and the ‘black box’ nature of DL models hinders efficient diagnosis and repair when they fail. Consequently, it is not safe to deploy DL models into safety-critical domains [3, 4]. Earlier studies proposed two lines of approaches to address this challenge. The first line of studies developed interpretability/explainability algorithms that shed light on operating principles underlying their decision-making [5, 6]. While the details may vary, they correlate DL models’ answers with input attributes or hidden layer representations and evaluate the importance of individual features on models’ decisions [7, 8, 9, 10, 11, 12]. The second line of studies sought algorithms than can evaluate the robustness of DL’s decision-making and enhance it [13, 14, 15, 16].
@@
We note that the first stage of selective attention can correspond to DL models’ forward passes and that the second stage requires high-level prior knowledge (e.g., visual features of police vehicles), which may not be readily available for DL models. We propose that foundation models trained on web-scale datasets may serve as proxy reservoirs of prior knowledge essential for top-down processes. We are particularly interested in the two segmentation models DINO [19] and Segment Anything model (SAM) [20], which detect a wide range of objects, since they can work with text-prompts; see also [21]. Specifically, we obtain an original prediction from a DL model and use it as a text prompt to identify the region of interest (ROI), which is used to confirm or reject the original prediction.
This certification process is referred to as ‘second-thought certification (STCert)’. First, we test if STCert can make DL models’ decision process more robust. Then, we test if STCert can alleviate DL models’ vulnerabilities to adversarial examples. Our empirical evaluations suggest STCert can enhance the robustness of DL and that STCert can warn adversarially manipulated inputs at high accuracy and natural adversarial examples 50% of time.
@@ 2 Approach
STCert mimics the two stage process of selective attention (Fig. 1). In the first stage, we obtain the original prediction from a DL model, which is an alien in the figure. In the second stage, we use two foundation models (DINO [20] and SAM [19]) to detect a ROI related to the original prediction (i.e., alien in Fig. 1) and use it as a secondary input to obtain a second thoughts prediction. In this study, we use an open-source package named ‘lang-SAM’ [21] to combine these two models. If a single ROI is identified, we test whether the original and second thought predictions are the same or not. If multiple ROIs are identified, we test whether the original answer is the same as one of the second thought predictions. If the original prediction is the same as or a part of second thought predictions, we consider the original answer a certified prediction. Below, we examine two different types of STCert, which are naive STCert and context-aware STCert.
@@ 2.1 Normal Dataset
We note that the ImageNet hierarchy (inherited from WordNet) is not homogeneous for all classes. For instance, ‘German Shepherd’ has 6 hypernyms, whereas ‘Jay’ class has 8 hypernyms, making the comparison of classes’ categories difficult. To address this issue, we adopt the subsets of ImageNets curated to study DL’s robustness [24]. Specifically, we use predetermined 5 subsets, Mixed_10, Mixed_13, Living_9, Big_12 and geirhos_16. These subsets (referred to as datasets below) contain ImageNet classes belonging to 10, 13, 9, 12 and 16 categories (a.k.a. superclasses), and we use the robustness package to select validation images from the selected classes/categories. In this study, we use validation examples included in these subsets to evaluate STCert. We do not modify or fine-tune DL models, and thus, DL models can choose one of 1,000 fine-grained classes as a prediction. We use NLTK toolkit [25] to identify the prediction’s hypernyms and determine if the prediction is hyponyms of any categories in the datasets; we ignore other hypernyms different from the categories in the database.
@@ 2.2 Adversarial inputs
We further evaluate whether STCert can alleviate DL models’ vulnerabilities to adversarial inputs. In this study, we consider both artificial and natural adversarial inputs. Artificial adversarial inputs are created by using projected gradient descent (PGD) attack [26], a white-box gradient-based attack, since PGD is one of the strongest adversarial attacks. PGD iterative updates following Eq. 1.
, where and denotes images and labels, , is a loss function (Cross-Entropy in our study), is a DL model. The clip function projects the perturbation to be less than or equal to 0.01 in our experiment. We use Advertorch [27], an open-source adversarial toolbox, to implement PGD attacks. For natural adversarial inputs, we use ImageNet-A [28], which contains naturally occurring examples that confuse a wide range of ImageNet models.
@@ 2.3 ImageNet models
We evaluate STCert using a set of ImageNet models, ResNet18, ResNet50 [29], VGG19_bn (with batch normalization) [30], DenseNet121 [31], ResNext101 [32] and ViT [33]. All models are implemented via timms python package [34] and selected to produce diverse architectures and complexity. It should be noted that all models are not modified or fine-tuned on the datasets (i.e., Imagenet subsets) discussed above.
@@ 2.4 Naive STCert
In the first and naive approach, we first create a mask covering ROI associated with an original prediction by combining DINO and SAM (see [21] for more details) and use it to remove all pixels outside the identified ROI (Eq. 2). Once all non-relevant pixels are removed, we use a common ImageNet preprocessing to create inputs for the second thought predictions.
,where , and denote an image, binary mask and blank image; where denote the element-wise product.
@@
Next, we compare certified predictions’ accuracy with that of original predictions. In the experiment, we split certified predictions into three categories, correct answers (CertCorr), inter-category error (InterError) and intra-category error (IntraError). We also observe how often STCert would reject the correct original predictions, which is indicated as ‘miss’ trials. As shown in Fig. 2, the accuracy of certified predictions range from 0.65 to 0.74 depending on the models. The missing and InterError rates are much lower than IntraError. To determine whether certified predictions are more accurate and more reliable than original predictions, we compare IntraError and InterError between original and certified predictions. The comparison in Fig. 2 shows that the error rates are substantially reduced, especially in InterError. We observe the same trend in the other 4 datasets (Fig. 3).
Further, we visually inspect the inter-category errors in the certified answers in Mixed_10. In 6 cases (Supplementary Fig. 1A-F), we find the ground truths and the certified predictions co-exist in the images. In some cases, the incorrect certified predictions and the ground truth labels are closely related. For instance, a certified prediction is ‘paddle’, when the ground truth is a canoe (Supplementary Fig. 1G). In 1 case (Supplementary Fig. 1H), ground truth labels are in fact incorrect.
@@ 2.5 Context-aware STCert
In the naive approach above, all pixels unrelated to the object originally predicted are zeroed out, which makes the inputs have substantially different statistics (e.g., mean pixel values). Since DL models can learn to use pixel statistics, we assume that these zeroed out pixels may disrupt DL models’ operations too strongly and that filling them out with natural pixels may restore the accuracy.
To obtain a second thought predictions determined by the original prediction only, the ideal backgrounds need to be related to the original prediction or semantically neutral. As it is not clear how to synthesize semantically neutral backgrounds, we resize ROIs to the input sizes native to the DL models Fig. 1. Specifically, we infer the bounding box of the object originally predicted, crop it out and resize it to the 224-by-224 (native input size to ImageNet models). With these ROIs with additional backgrounds, we evaluate the rates of CertCorr, IntraError, InterError, Miss and No-box trials and report them in Fig. 4A. As shown in the figure, the CertCorr rate is increased for all 6 models. We further compare InerError and IntraError between original and certified predictions and made two observations (Fig. 4B). First, both IntraError and InterError rates are reduced. Second, InterError is reduced more significantly than IntraError. We find equivalent results in all other datasets (Fig. 5).
Further, inspired by the observation that DL models utilize contexts to make decisions, we gradually add more background pixels to the extracted features. As the backgrounds surrounding ROIs contain contexts closely related to original predictions, we enlarge bounding boxes. To evaluate the influence of these backgrounds (i.e., contexts), we test 5 different bounding box sizes that are referred to as context width (cw) in Fig. 6. As seen in Fig. 7A, the second thought predictions’ accuracy increases, as the bounding boxes grow in terms of sizes. More importantly, the growth of certified predictions’ errors remain lower than the increase of certified predictions’ accuracy, suggesting that adding backgrounds improves the accuracy of certified predictions in exchange for limited increase in errors. We repeat the same experiment with other datasets and find the equivalent results (Fig. 7B-E).
We observe that certified predictions and ground truths are often semantically related to each other, even when they belong to other categories. We quantify this observation by estimating distances between the ground truth labels and the certified predictions using spaCy similarity measure [35]. Supplemental Table 1 shows the most similar pairs of ground truth labels and certified predictions.
@@ 2.6 Implications for the relationships between contexts and models’ decision-making
STCert use context to make certified decisions, and the amount of context can be controlled with context width (CW, Fig. 6). If second thought predictions and original predictions of a DL model are always the same without context, we can interpret this result as evidence suggesting that the model does not utilize context for its decision. By contrast, if second thought predictions and original predictions become similar only when a large amount of context is provided, we can interpret it as evidence supporting that context is crucial for the model’s decision. With this possibility in mind, we evaluate models’ sensitivity to the context amount by evaluating the certified prediction accuracy depending on cw (Fig. 8) and find that all models’ benefit from context but at different levels. VITs accuracy increases most strongly, and ResNet50 appears to be least sensitives to the context. Interestingly, although ResNet18 and ResNet50 share the same architecture, ResNet18 requires more context to make proper decisions. Conversely, ResNet50 may utilize information encoded in ROIs (i.e., internal features) more effectively. Together with that ResNet50 has a deeper architecture, we speculate that the depth may enhance the model’s capability to learn internal features.
Then, we ask if models’ sensitivity to contexts depend on the identity of foreground objects. To address this question, we further analyze the accuracy of certified predictions depending on the ground truth labels and find that the models’ sensitivity strongly depends on the class (Fig. 9). Interestingly, we note that all 6 models strongly benefited from context strongly, when the ground truth label is a ‘canoe’. As canoes have simple internal features, contexts may be essential for DL models to tell it from other objects. Similarly, the certified predictions’ accuracy improved strongly when DL models make decisions on other objects, which have internal features common for multiple fine-grained classes under the same superclass; see Table 2 for 10 objects, on which context improve the accuracy of certified predictions. These results raise the possibility that DL models may learn to use contexts when internal features are not sufficient for making proper decisions.
If we better understand what controls DL models’ sensitivity to contexts (and objects’ internal features), we could select the optimal architecture depending on the dataset properties related to the complexity of internal features. We believe STCert is a good starting point to develop analysis tools that evaluates DL models’ sensitivity to contexts.
@@ 3.1 Artificial adversarial examples
To test the utility of STCert, we ask if second thought predictions are the same as original predictions to perturbed inputs. If they are different, STCert rejects original answers and tag them as unreliable predictions, which is desirable. In this study, we use a PGD untargeted attack to create adversarial examples from all validation images in the Mixed_10. We use an untartgeted attack since it is the simpler attack than targeted one and can build the stronger attack. Using this untargeted attack, we aim to establish the lower bound of STCert’ s capability of detecting DL models’ disrupted/abnormal decision process.
The created images are highly effective and roughly 95% images can fool DL models. We apply STCert to incorrect predictions to these perturbed inputs, which means that an original prediction (i.e., a text-prompt) instructs a segmentation model to search for an object, which likely does not exist in an input. The segmentation models do not find any relevant object more often (Fig. 10), and the rates of certified predictions are extremely low for all the tested ImageNet classifiers (see green bars in Fig. 10), suggesting that STCert can reliably reject adversarially manipulated predictions/decisions.
@@ 3.2 Natural adversarial examples
Notably, DL models’ vulnerability is not limited to artificially-created adversarial inputs. Some naturally occurring visual patterns can be extremely difficult to detect for DL models. Some natural adversarial examples may contain benign objects, but they can effectively confuse and fool a wide range of DL models, which may prevent them from being deployed in real-world applications. In our study, we ask if STCert can reject DL models’ decisions on natural adversarial examples using ImageNet-A, which includes multiple images from 200 ImageNet classes. We make two main observations. First, the segmentation model cannot find the relevant objects in approximately 10% of the examples (Fig. 11). Second, close to 50% of the inputs are not certified (Fig. 11). That is, our certification process can reject original predictions on natural adversarial examples 50% of the time.
These results suggest that STCert can detect anomalous decisions induced by artificial or natural adversarial examples, helping DL models make more reliable decisions. Then, why does the certification work far better on artificial adversarial inputs than on natural adversarial inputs? The adversarial perturbations, most strongly, change the internal representations not directly related to the foreground objects (i.e., objects determining images’ labels) [38], which raises the possibility that adversarial perturbations can allow pixels in backgrounds to exert dominant influence on DL models’ decision-making process. By design, STCert can natively remove backgrounds’ influence and thus can detect adversarial examples.
Further, although our understanding of DL models’ difficulty with natural adversarial examples remains poor, we note that natural adversarial objects have somewhat complex textures, opening up multiple interpretations. Even when STCert isolates the objects from the backgrounds, the complex textures of the objects can still deceive STCert. If the complex textures lead to natural adversarial inputs, a natural extension of STCert may improve the accuracy of STCert on adversarial inputs. In the current STCert, we use the same model for both original and second thought predictions, but instead of using the same model twice, we could train a secondary model for the second thought predictions to be less sensitive to the textures of the objects.

### v2

@@ 1 Introduction
Deep learning (DL) has excelled in various domains including computer vision and natural language processing [1, 2]. As deep neural networks or DL models rely on a large amount of linear/nonlinear operations, their operations are intractable, and the ‘black box’ nature of DL models hinders efficient diagnosis and repair when they fail. Consequently, it is not safe to deploy DL models into safety-critical domains [3, 4]. Earlier studies proposed two lines of approaches to address this challenge. The first line of studies developed interpretability/explainability algorithms that shed light on operating principles underlying their decision-making [5, 6]. While the details may vary, they correlate DL models’ answers with input attributes or hidden layer representations and evaluate the importance of individual features on models’ decisions [7, 8, 9, 10, 11, 12]. The second line of studies sought algorithms that can evaluate the robustness of DL’s decision-making and enhance it [13, 14, 15, 16].
@@
The first stage of selective attention can correspond to DL models’ forward passes, and the second stage requires high-level prior knowledge (e.g., visual features of police vehicles), which may not be readily available for DL models. We propose that foundation models trained on web-scale datasets may serve as proxy reservoirs of prior knowledge essential for top-down processes. We are particularly interested in foundational segmentation models, DINO [19], Grounding DINO [20] and Segment Anything model (SAM) [21], which can detect a wide range of objects and work with diverse prompts; see also [22]. Specifically, we obtain an original prediction from a DL model and use it as a text prompt to identify the region of interest (ROI), which is used to confirm or reject the original prediction.
This certification process is referred to as ‘second-thought certification (STCert)’. First, we test if STCert can make DL models’ decision process more robust. Then, we test if STCert can alleviate DL models’ vulnerabilities to adversarial examples. Our empirical evaluations suggest that STCert can enhance the robustness of DL and that STCert can warn adversarially manipulated inputs at high accuracy and natural adversarial examples 50% of time.
@@ 2 Approach
STCert mimics the two stage process of selective attention (Fig. 1). In the first stage, we obtain the original prediction from a DL model, which is an alien in the figure. In the second stage, we use two foundation models (Grounding DINO [20] and SAM [21]) to detect a ROI related to the original prediction (i.e., alien in Fig. 1) and use it as a secondary input to obtain a second thoughts prediction. In this study, we use an open-source package named ‘lang-SAM’ [22] to combine these two models. If a single ROI is identified, we test whether the original and second thought predictions are the same or not. If multiple ROIs are identified, we test whether the original answer is the same as one of the second thought predictions. If the original prediction is the same as or a part of second thought predictions, we consider the original answer a certified prediction. If no ROI is identified with the fine-grained class name, its super-class is used as an alternative prompt. Below, we examine two different types of STCert, which are naive STCert and context-aware STCert.
@@ 2.1 Normal Dataset
We note that the ImageNet hierarchy (inherited from WordNet) is not homogeneous for all classes. For instance, ‘German Shepherd’ has 6 hypernyms, whereas ‘Jay’ class has 8 hypernyms, making the comparison of classes’ categories difficult. To address this issue, we adopt the subsets of ImageNets curated to study DL’s robustness [25]. Specifically, we use predetermined 5 subsets, Mixed_10, Mixed_13, Living_9, Big_12 and geirhos_16. These subsets (referred to as datasets below) contain ImageNet classes belonging to 10, 13, 9, 12 and 16 categories (a.k.a. super-classes), and we use the robustness package [25] to select validation images from the selected classes/categories. In this study, we use validation examples included in these subsets to evaluate STCert. We do not modify or fine-tune DL models, and thus, DL models can choose one of 1,000 fine-grained classes as a prediction. We use NLTK toolkit [26] to identify the prediction’s hypernyms and determine if the prediction is hyponyms of any categories in the datasets; we ignore other hypernyms different from the categories in the database.
@@ 2.2 Adversarial inputs
We further evaluate whether STCert can alleviate DL models’ vulnerabilities to adversarial inputs. In this study, we consider both artificial and natural adversarial inputs. Artificial adversarial inputs are created by using autoattack [27], which relies on two custom Projected Gradient Descent (PGD), FAB [28] and Square Attack [29]. We use standard autoattack with and norm=‘Linf’. For natural adversarial inputs, we use ImageNet-A [30], which contains naturally occurring examples that confuse a wide range of ImageNet models.
@@ 2.3 ImageNet models
We evaluate STCert using a set of ImageNet models, ResNet18, ResNet50 [31], VGG19_bn (with batch normalization) [32], DenseNet121 [33], ResNext101 [34] and vision transformer (VIT) [35]. All models are implemented via Timms python package [36]. It should be noted that all models are not modified or fine-tuned on the datasets (i.e., Imagenet subsets) discussed above. All experiments and analysis are conducted using a workstation with Intel core 9 CPU and RTX 4090.
@@ 2.4 Naive STCert
In the first and naive approach, we first create a mask covering ROI associated with an original prediction by combining Grounding DINO and SAM (see [22] for more details) and use it to remove all pixels outside the identified ROI (Eq. 1). Once all non-relevant pixels are removed, we use a common ImageNet preprocessing to create inputs for the second thought predictions.
,where , and denote an image, binary mask and blank image; where denotes the element-wise product.
@@
Next, we compare certified predictions’ accuracy with that of original predictions. In the experiment, we split certified predictions into three categories, correct answers (CertCorr), inter-category error (InterError) and intra-category error (IntraError). We also observe how often STCert would reject the correct original predictions, which is indicated as ‘miss’ trials. As shown in the top row of Fig. 2A, the accuracy of certified predictions range from 0.65 to 0.74 depending on the models. The miss and InterError rates are much lower than IntraError. To determine whether certified predictions are more accurate and more reliable than original predictions, we compare IntraError and InterError between original and certified predictions. As shown in the bottom row of Fig. 2A, the error rates are substantially reduced, especially in InterError. We observe the same trend in the other 4 datasets (Supplementary Fig. 1).
Further, we visually inspect the inter-category errors in the certified answers in Mixed_10. In 6 cases (Supplementary Fig.2 A-F), we find the ground truths and the certified predictions co-exist in the images. In some cases, the incorrect certified predictions and the ground truth labels are closely related. For instance, a certified prediction is ‘paddle’, when the ground truth is a canoe (Supplementary Fig. 2). In 1 case (Supplementary Fig. 2H), ground truth labels are in fact incorrect.
@@ 2.5 Context-aware STCert
In the naive approach above, all pixels unrelated to the object originally predicted are zeroed out, which makes the inputs have substantially different statistics (e.g., mean pixel values). Since DL models can learn to use pixel statistics, we assume that these zeroed out pixels may disrupt DL models’ operations too strongly and that filling them out with natural pixels may restore the accuracy. To obtain a second thought predictions determined by the original prediction only, the ideal backgrounds need to be related to the original prediction or semantically neutral. As it is not clear how to synthesize semantically neutral backgrounds, we resize ROIs to the input sizes native to the DL models ( Fig. 1). Specifically, we infer the bounding box of the object originally predicted, crop it out and resize it to the 224-by-224 (native input size to ImageNet models). With these ROIs with additional backgrounds, we evaluate the rates of CertCorr, IntraError, InterError, Miss and No-box trials and report them in the top row of Fig. 2B. As shown in the figure, the CertCorr rate is increased for all 6 models. We further compare InterError and IntraError between original and certified predictions and made two observations (Bottom row of Fig. 2B). First, both IntraError and InterError rates are reduced. Second, InterError is reduced more significantly than IntraError. We find equivalent results in all other datasets (Supplementary Fig. 3).
Further, inspired by the observation that DL models utilize contexts to make decisions, we gradually add more background pixels to the extracted features. As the backgrounds surrounding ROIs contain contexts closely related to original predictions, we enlarge bounding boxes. To evaluate the influence of these backgrounds (i.e., contexts), we test 5 different bounding box sizes that are referred to as context width (cw) in Fig. 3A. As seen in Fig. 3B, the second thought predictions’ accuracy increases, as the bounding boxes grow in terms of sizes. More importantly, the growth of certified predictions’ errors remain lower than the increase of certified predictions’ accuracy, suggesting that adding backgrounds improves the accuracy of certified predictions in exchange for limited increase in errors. We repeat the same experiment with other datasets and find the equivalent results (Fig. 3C-F).
We observe that certified predictions and ground truths are often semantically related to each other, even when they belong to other categories. We quantify this observation by estimating distances between the ground truth labels and the certified predictions using spaCy similarity measure [37]. Supplementary Table 1 shows the most similar pairs of ground truth labels and certified predictions.
@@ 2.6 Implications for the relationships between contexts and models’ decision-making
STCert uses context to make certified decisions, and the amount of context can be controlled with context width (CW, Fig. 3A). If second thought predictions and original predictions of a DL model are always the same without context, we can interpret this result as evidence suggesting that the model does not utilize context for its decision. By contrast, if second thought predictions and original predictions become similar only when a large amount of context is provided, we can interpret it as evidence supporting that context is crucial for the model’s decision. With this possibility in mind, we evaluate models’ sensitivity to the context amount by evaluating the certified prediction accuracy depending on cw (Fig. 4) and find that all models benefit from context but at different levels. VIT’s accuracy increases most strongly, and ResNet50 appears to be least sensitive to the context. Interestingly, although ResNet18 and ResNet50 share the same architecture, ResNet18 requires more context to make proper decisions. Conversely, ResNet50 may utilize information encoded in ROIs (i.e., internal features) more effectively. Together with that ResNet50 has a deeper architecture, we speculate that the depth may enhance the model’s capability to learn internal features.
We note that models’ sensitivity to contexts depend on the identity of foreground objects (Supplementary Fig. 4). Interestingly, all 6 models strongly benefit from context, when the ground truth label is a ‘canoe’. As canoes have simple internal features, contexts may be essential for DL models to tell it from other objects. Similarly, the certified predictions’ accuracy improved strongly when DL models make decisions on other objects, which have internal features common for multiple fine-grained classes under the same super-class; see Supplementary Table 2 for 10 objects, on which contexts improve the accuracy of certified predictions. These results raise the possibility that DL models may learn to use contexts when internal features are not sufficient for making proper decisions. If we better understand what controls DL models’ sensitivity to contexts (and objects’ internal features), we could select the optimal architecture depending on the dataset properties related to the complexity of internal features. We believe STCert is a good starting point to develop analysis tools that evaluate DL models’ sensitivity to contexts.
@@ 3.1 Artificial adversarial examples
We further ask if second thought predictions are the same as original predictions to perturbed inputs. If they are different, STCert would reject original answers and tag them as unreliable predictions, which is desirable. In this study, we use ‘Auto Attack’, one of the strongest attacks [27], to create adversarial examples from all validation images in Mixed_10. The adversarial perturbations crafted by Auto Attack are highly effective, leading all 6 DL models to make incorrect predictions on all crafted adversarial inputs; that is, all 6 model’s accuracies go down to 0%. As we aim to test STCert’s ability to reject adversarially manipulated (incorrect) predictions, we forward the adversarially manipulated predictions and adversarial inputs to STCert and measure how many times STCert would reject or confirm predictions. When STCert rejects predictions, STCert successfully detects adversarial perturbation. When STCert confirms predictions, STCert fails to detect adversarial perturbation. We also measure the number of times STCert fails to detect the bounding boxes (ROI) of predicted objects, as this failure can also indicate original predictions that should be inspected further (i.e., successful detection of adversarial inputs).
Our results suggest that STCert detection rates strongly depend on the model (Fig. 5A). The success rates are 22% for VGG19 and 87.5% for VIT. For other models, it is roughly 40-50%. This relatively low detection rate may be explained by the fact that adversarially manipulated predictions and normal predictions often belong to the same category, and thus, the segmentation model can still find proper ROIs. With the correct ROI, the second classification can also be highly vulnerable to the adversarial inputs, since the original and second thought predictions rely on the same classifier, for which the adversarial perturbations are crafted. Thus, we can hypothesize that when original and second thought predictions are made by two different models, STCert can detect adversarial inputs more accurately. For instance, ResNet18 is used for the original prediction, but ResNet50 is used for the second thought prediction.
To address this hypothesis, we use two different classifiers for STCert and find the detection rates to be much higher and close to 70-80% (Fig. 5B), when a combination of ResNet18 and ResNet50 are used. We further test all possible combinations for 6 classifiers () and measure how often STCert detects adversarial inputs and fails to detect bounding boxes. That is, we add the blue and orange bars in (Fig. 5) to estimate the ‘overall success rate’. These overall success rates are 70% or higher for all combinations (Table 2), raising the possibility that STCert, which utilizes two equivalent but distinct models, can reliably detect adversarial inputs.
R18 | R50 | V19 | D121 | RN | VIT | |
R18 | 0.37 | 0.80 | 0.72 | 0.74 | 0.81 | 0.86 |
R50 | 0.73 | 0.49 | 0.72 | 0.73 | 0.77 | 0.83 |
V19 | 0.77 | 0.83 | 0.22 | 0.79 | 0.81 | 0.88 |
D121 | 0.81 | 0.85 | 0.81 | 0.51 | 0.85 | 0.89 |
RN | 0.75 | 0.75 | 0.72 | 0.71 | 0.34 | 0.83 |
VIT | 0.91 | 0.91 | 0.91 | 0.91 | 0.92 | 0.88 |
@@ 3.2 Natural adversarial examples
Notably, some naturally occurring visual patterns can be extremely difficult to detect for DL models. These natural adversarial examples may contain benign objects, but they can effectively confuse and fool a wide range of DL models, which may prevent them from being deployed in real-world applications. In our study, we ask if STCert can reject DL models’ decisions on natural adversarial examples using ImageNet-A, which includes multiple images from 200 ImageNet classes. As in artificial adversarial perturbations, we test overall success rate (i.e., combination of detection success trials and no-box trials) using all possible pairs of first and second classifiers (). Once again, the first classifier makes original predictions, and the second classifier confirms/rejects the original predictions. As shown in Table 3, the overall success rate is higher when the first and second classifiers are different.
R18 | R50 | V19 | D121 | RN | VIT | |
R18 | 0.49 | 0.68 | 0.65 | 0.64 | 0.69 | 0.75 |
R50 | 0.51 | 0.49 | 0.67 | 0.66 | 0.71 | 0.77 |
V19 | 0.50 | 0.70 | 0.22 | 0.49 | 0.81 | 0.66 |
D121 | 0.51 | 0.70 | 0.67 | 0.49 | 0.70 | 0.77 |
RN | 0.52 | 0.70 | 0.67 | 0.66 | 0.49 | 0.77 |
VIT | 0.53 | 0.71 | 0.68 | 0.67 | 0.72 | 0.49 |
3.3 Two classifier-based STCert on normal inputs
The results above suggest that STCert can detect incorrect predictions on artificial and natural adversarial examples more reliably when two different classifiers are used for STCert. Then, does STCert work better in general when two classifiers are used to make original and second thought predictions? To address this question, we test two classifier-based ST on normal test examples in Mixed_10. We observe that STCert can still 1) confirm correct answers (Fig. 6), and 2) it detects incorrect answers more reliably ( Fig. 2 VS Fig. 6), when two different classifiers are used for STCert. As using two classifiers may increase the necessary cost for STCert, the two classifier-based STCert may be suitable for high-stakes and risk problems, which can justify the cost.